# The theory of implicit operations

### v1

@@ 11. The Beth companion
When $\mathsf{K}=\mathsf{M}$, part (i) of the above definition specializes to the familiar demand that the implicit operation $f$ be interpolated by the set of terms $\{t_{i}:i\in I\}$. Notice that part (ii) subsumes part (i), for term functions can be viewed as implicit operations (see 3.8). However, we opted for splitting the definition in two halves for the sake of clarity. Lastly, we remark that the above definition applies to the situation where $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ because in this case ${L}_{\mathsf{K}}\subseteq{L}_{\mathsf{M}}$ and the ${L}_{\mathsf{K}}$-reducts of $\mathsf{M}$ belong to $\mathsf{K}$ (see Proposition 10.2). ∎
@@
Let $\mathsf{K}$ be a universal class, $\mathcal{F}=\mathsf{ext}_{pp}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Then $\mathsf{K}$ has a Beth companion if and only if $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a Beth companion of $\mathsf{K}$.
@@
    every member of $\mathsf{imp}_{pp}(\mathsf{K})$ is interpolated in $\mathsf{M}$ by a set of terms of $\mathsf{M}$.
@@
    every member of $\mathsf{imp}_{pp}(\mathsf{K})$ is interpolated in $\mathsf{M}$ by a single term of $\mathsf{M}$.
@@
(ii)–(v): Analogous to the proof of (i). For (ii), use Theorems 10.12(ii) and 7.5. For (iii), use Theorems 10.12(i) and 7.5. For (iv), use 10.17 and the fact that the variety of implicative semilattices has the strong epimorphism surjectivity property (see [42, Props. 81 & 82]). For (v), use 10.19 and the fact that the variety of Heyting algebras of depth $\leqslant 2$ has the strong epimorphism surjectivity property (see [91, Thm. 8.1(3)]).
(vi): Consider a universal class $\mathsf{K}$ with the strong epimorphism surjectivity property. In view of Example 10.8, the class $\mathsf{K}$ is a pp expansion of itself. Therefore, we can apply 11.6, obtaining that $\mathsf{K}$ is a Beth companion of itself. ∎
@@
The following shows that the hypothesis that $\mathsf{K}$ is a quasivariety in 11.7 cannot be replaced with the requirement that $\mathsf{K}$ is a universal class. Let $\bm{A}=\langle A;\land,\lor,0,1\rangle$ be the two-element bounded lattice and $\mathsf{K}=\mathbb{U}(\bm{A})$. Since $\bm{A}$ is finite and has no proper subalgebras, from 2.2 and 2.14 it follows that $\mathsf{K}=\mathbb{U}(\bm{A})=\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\bm{A})=\mathbb{I}(\bm{A})$. Therefore, every member of $\mathsf{K}$ lacks proper subalgebras because it is isomorphic to $\bm{A}$. Therefore, for all $\bm{B}\leqslant\bm{C}\in\mathsf{K}$ we have $\mathsf{d}_{\mathsf{K}}(\bm{B},\bm{C})=\mathsf{d}_{\mathsf{K}}(\bm{C},\bm{C})=\bm{C}$. Together with 6.6, this yields that $\mathsf{K}$ has the strong epimorphism surjectivity property. Therefore, 11.9(vi) implies that $\mathsf{K}$ is a Beth companion of itself.
We will show that there exists another Beth companion $\mathsf{M}$ of $\mathsf{K}$ such that $\mathsf{K}$ and $\mathsf{M}$ are not faithfully term equivalent relative to $\mathsf{K}$. Since $\mathsf{K}$ is a class of bounded distributive lattices, it follows from 3.20(ii) that taking complements yields a unary implicit operation $f$ of $\mathsf{K}$ defined by a conjunction of equations. As $f^{\bm{B}}$ is total for every $\bm{B}\in\mathbb{I}(\bm{A})=\mathsf{K}$, we have $f\in\mathsf{ext}_{eq}(\mathsf{K})$. Let ${L}_{f}={L}_{\mathsf{K}}\cup\{\neg\}$ be an $f$-expansion of ${L}_{\mathsf{K}}$ and $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{f}])$ the pp expansion of $\mathsf{K}$ induced by $f$ and ${L}_{f}$. As every member of $\mathsf{K}$ lacks proper subalgebras, so does every member of $\mathsf{K}[{L}_{f}]$, and hence $\mathsf{M}=\mathsf{K}[{L}_{f}]$. Together with $\mathsf{K}=\mathbb{I}(\bm{A})$ and the fact that $\bm{A}[{L}_{f}]$ is defined (because $f^{\bm{A}}$ is total), this yields $\mathsf{M}=\mathbb{I}(\bm{A}[{L}_{f}])$. By arguing as above we obtain that $\mathsf{M}$ has the strong epimorphism surjectivity property. Therefore, 11.6 implies that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$.
@@
Recall from the assumptions that $\mathsf{M}_{1}$ is a pp expansion of $\mathsf{K}$ and that $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$. Therefore, $\mathsf{M}_{2}$ is also a pp expansion of $\mathsf{K}$ by 10.6. This fact will be used repeatedly in the proof.
@@
We will prove that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\in\mathsf{M}_{1}$. Recall from the assumptions that $\mathsf{K}$ is a universal class. Therefore, so is its pp expansion $\mathsf{M}_{1}$ by 10.3(i). Together with the assumption that $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$, this allows us to apply 9.6, obtaining $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\in\mathsf{M}_{1}$, as desired. Together with the above display and the assumption that $\{t_{i}:i\in I\}$ interpolates $f$ in $\mathsf{M}_{1}$, this implies that there exists $i\in I$ such that $f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})=t_{i}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}(a_{1},\dots,a_{n})=t_{i}^{\bm{A}}(a_{1},\dots,a_{n})$.
(ii): Assume that $\mathsf{M}_{2}$ is a Beth companion of $\mathsf{M}_{1}$. We will prove that $\mathsf{M}_{2}$ is a Beth companion of $\mathsf{K}$ as well. As in the previous case, it suffices to show that every implicit operation of $\mathsf{K}$ is interpolated in $\mathsf{M}_{2}$ by a family of terms of $\mathsf{M}_{2}$. Accordingly, consider an $n$-ary $f\in\mathsf{imp}(\mathsf{K})$. By 10.26 there exists $g\in\mathsf{imp}(\mathsf{M}_{1})$ such that $g^{\bm{A}}=f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{M}_{1}$. Observe that $g$ is interpolated in $\mathsf{M}_{2}$ by a family $\{t_{i}:i\in I\}$ of terms of $\mathsf{M}_{2}$ because $\mathsf{M}_{2}$ is a Beth companion of $\mathsf{M}_{1}$. We will prove that the same family interpolates $f$ in $\mathsf{M}_{2}$. To this end, consider $\bm{A}\in\mathsf{M}_{2}$ and $a_{1},\dots,a_{n}\in A$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}})$. Since $\mathsf{K}$ is a universal class, 10.3(i) implies that $\mathsf{M}_{1}$ is a universal class as well. As $\bm{A}\in\mathsf{M}_{2}$ and $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$ we have $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\in\mathsf{M}_{1}$ by Proposition 9.6. Furthermore, $(\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}){\upharpoonright}_{{L}_{\mathsf{K}}}=\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$ because ${L}_{\mathsf{K}}\subseteq{L}_{\mathsf{M}_{1}}$. Therefore,
@@
Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, we may assume that it is induced by some $\mathcal{G}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and $\mathcal{G}$-expansion ${L}_{\mathcal{G}}$ of ${L}_{\mathsf{K}}$. As $\mathcal{F}=\mathsf{ext}_{pp}(\mathsf{K})$ by assumption, we have $\mathcal{G}\subseteq\mathsf{ext}_{pp}(\mathsf{K})=\mathcal{F}$. Then there exists an $\mathcal{F}$-expansion ${L}_{\mathcal{F}}^{\prime}$ of the form ${L}_{\mathcal{G}}\cup\{g_{f}:f\in\mathcal{F}-\mathcal{G}\}$. Since $\mathcal{G}\subseteq\mathcal{F}$ and ${L}_{\mathcal{G}}\subseteq{L}_{\mathcal{F}}^{\prime}$, 10.7 implies that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}^{\prime}])$ is a pp expansion of $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$. Together with the assumption that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, this allows us to apply Theorem 11.4(i), obtaining that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}^{\prime}])$ is a Beth companion of $\mathsf{K}$ as well. Lastly, the definition of ${L}_{\mathcal{F}}^{\prime}$ guarantees that the classes $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ and $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}^{\prime}])$ are term equivalent. ∎
@@
Applying (iv) to $g\in\mathsf{imp}_{pp}(\mathsf{K})$ yields a set $\{t_{i}:i\in I\}$ of terms of $\mathsf{M}$ that interpolates $g$ in $\mathsf{M}$. We will show that $\{t_{i}:i\in I\}$ interpolates $f$ as well. To this end, consider $\bm{A}\in\mathsf{M}$ and $a_{1},\dots,a_{n}\in A$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$. Since $\bm{A}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, we have $\bm{A}\leqslant\bm{B}$ for some $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$. As $f\in\mathsf{imp}(\mathsf{M})$ and $\bm{A},\bm{B}\in\mathsf{M}$ are such that $\bm{A}\leqslant\bm{B}$, from 8.1 and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$ it follows that
@@
It only remains to prove 11.7. The proof hinges on the next observation.
@@
Let $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ be a pair of Beth companions of a quasivariety $\mathsf{K}$. We will prove that $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are faithfully term equivalent relative to $\mathsf{K}$. For $i=1,2$ let $\mathcal{F}_{i}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ be such that $\mathsf{M}_{i}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{i}}])$. We may assume that
@@
Then let $T_{i}$ be the set of terms of ${L}_{\mathcal{F}_{i}}$ with variables in $\{x_{n}:n\in\mathbb{N}\}$. Since $\mathsf{K}$ is a quasivariety, 11.6 implies that for every $n$-ary $f\in\mathsf{imp}_{pp}(\mathsf{K})$ and $i=1,2$ there exists an $n$-ary term $F_{i}(f)\in T_{i}$ that interpolates $f$ in $\mathsf{M}_{i}$. Therefore, for every $n$-ary $f\in\mathsf{imp}_{pp}(\mathsf{K})$, $\bm{A}\in\mathsf{M}_{i}$ with $i=1,2$, and $a_{1},\dots,a_{n}\in A$,
@@
for each $f_{1}\in\mathcal{F}_{1}$ and $f_{2}\in\mathcal{F}_{2}$. We proceed to show that $\tau$ and $\rho$ witness that $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are faithfully term equivalent relative to $\mathsf{K}$. Let $\mathcal{F}=\mathcal{F}_{1}\cup\mathcal{F}_{2}$ and let ${L}_{\mathcal{F}}$ be an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ such that ${L}_{\mathcal{F}_{1}},{L}_{\mathcal{F}_{2}}\subseteq{L}_{\mathcal{F}}$. Recall that $\mathcal{F}=\mathcal{F}_{2}\cup\mathcal{F}_{2}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ by assumption. Therefore, 10.7 implies that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of both $\mathsf{M}_{1}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{1}}])$ and $\mathsf{M}_{2}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{2}}])$. To verify conditions (i) and (iii) in the definition of faithful term equivalence, consider $\bm{A}\in\mathsf{M}_{1}$. We need to prove that $\rho(\bm{A})\in\mathsf{M}_{2}$ and $\tau(\rho(\bm{A}))=\bm{A}$. Since $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{M}_{1}$, 10.2 implies that $\bm{A}$ is an ${L}_{\mathcal{F}_{1}}$-subreduct of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As every member of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a subalgebra of a member of $\mathsf{K}[{L}_{\mathcal{F}}]$, it follows that there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{1}}}$. Since $\mathsf{M}_{2}$ is a universal class by 10.3(i) and $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is also a pp expansion of $\mathsf{M}_{2}$, 10.2 implies that $\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{2}}}\in\mathsf{M}_{2}$.
We now show that $\rho(\bm{A})\in\mathsf{M}_{2}$. Observe that $\mathsf{M}_{2}$ is closed under subalgebras because it is a pp expansion of $\mathsf{K}$. Hence, it is sufficient to prove $\rho(\bm{A})\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{2}}}$ because $\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{2}}}\in\mathsf{M}_{2}$. Since $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{1}}}$, from 11.12(iii) it follows that
@@
Next we prove that $\tau(\rho(\bm{A}))=\bm{A}$. By 11.12(iii) we obtain

### v2

@@ 11. The Beth companion
When $\mathsf{K}=\mathsf{M}$, part (i) of the above definition specializes to the familiar demand that the implicit operation $f$ be interpolated by the set of terms $\{t_{i}:i\in I\}$. Notice that part (ii) subsumes part (i), for term functions can be viewed as implicit operations (see Example 3.8). However, we opted for splitting the definition in two halves for the sake of clarity. Lastly, we remark that the above definition applies to the situation where $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ because in this case ${L}_{\mathsf{K}}\subseteq{L}_{\mathsf{M}}$ and the ${L}_{\mathsf{K}}$-reducts of $\mathsf{M}$ belong to $\mathsf{K}$ (see Proposition 10.2). ∎
@@
Let $\mathsf{K}$ be a universal class, $\mathcal{F}=\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$, and ${L}_{\mathcal{F}}$ an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$. Then $\mathsf{K}$ has a Beth companion if and only if $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a Beth companion of $\mathsf{K}$.
@@
    every member of $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ is interpolated in $\mathsf{M}$ by a set of terms of $\mathsf{M}$.
@@
    every member of $\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ is interpolated in $\mathsf{M}$ by a single term of $\mathsf{M}$.
@@
(ii)–(v): Analogous to the proof of (i). For (ii), use Theorems 10.12(ii) and 7.5. For (iii), use Theorems 10.12(i) and 7.5. For (iv), use Theorem 10.17 and the fact that the variety of implicative semilattices has the strong epimorphism surjectivity property (see [42, Props. 81 & 82]). For (v), use Theorem 10.19 and the fact that the variety of Heyting algebras of depth $\leqslant 2$ has the strong epimorphism surjectivity property (see [91, Thm. 8.1(3)]).
(vi): Consider a universal class $\mathsf{K}$ with the strong epimorphism surjectivity property. In view of Example 10.8, the class $\mathsf{K}$ is a pp expansion of itself. Therefore, we can apply Theorem 11.6, obtaining that $\mathsf{K}$ is a Beth companion of itself. ∎
@@
The following shows that the hypothesis that $\mathsf{K}$ is a quasivariety in Theorem 11.7 cannot be replaced with the requirement that $\mathsf{K}$ is a universal class. Let $\bm{A}=\langle A;\land,\lor,0,1\rangle$ be the two-element bounded lattice and $\mathsf{K}=\mathbb{U}(\bm{A})$. Since $\bm{A}$ is finite and has no proper subalgebras, from Theorem 2.2 and Proposition 2.14 it follows that $\mathsf{K}=\mathbb{U}(\bm{A})=\mathbb{I}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\bm{A})=\mathbb{I}(\bm{A})$. Therefore, every member of $\mathsf{K}$ lacks proper subalgebras because it is isomorphic to $\bm{A}$. Therefore, for all $\bm{B}\leqslant\bm{C}\in\mathsf{K}$ we have $\mathsf{d}_{\mathsf{K}}(\bm{B},\bm{C})=\mathsf{d}_{\mathsf{K}}(\bm{C},\bm{C})=\bm{C}$. Together with Proposition 6.6, this yields that $\mathsf{K}$ has the strong epimorphism surjectivity property. Therefore, Theorem 11.9(vi) implies that $\mathsf{K}$ is a Beth companion of itself.
We will show that there exists another Beth companion $\mathsf{M}$ of $\mathsf{K}$ such that $\mathsf{K}$ and $\mathsf{M}$ are not faithfully term equivalent relative to $\mathsf{K}$. Since $\mathsf{K}$ is a class of bounded distributive lattices, it follows from Theorem 3.20(ii) that taking complements yields a unary implicit operation $f$ of $\mathsf{K}$ defined by a conjunction of equations. As $f^{\bm{B}}$ is total for every $\bm{B}\in\mathbb{I}(\bm{A})=\mathsf{K}$, we have $f\in\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$. Let ${L}_{f}={L}_{\mathsf{K}}\cup\{\neg\}$ be an $f$-expansion of ${L}_{\mathsf{K}}$ and $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{f}])$ the pp expansion of $\mathsf{K}$ induced by $f$ and ${L}_{f}$. As every member of $\mathsf{K}$ lacks proper subalgebras, so does every member of $\mathsf{K}[{L}_{f}]$, and hence $\mathsf{M}=\mathsf{K}[{L}_{f}]$. Together with $\mathsf{K}=\mathbb{I}(\bm{A})$ and the fact that $\bm{A}[{L}_{f}]$ is defined (because $f^{\bm{A}}$ is total), this yields $\mathsf{M}=\mathbb{I}(\bm{A}[{L}_{f}])$. By arguing as above we obtain that $\mathsf{M}$ has the strong epimorphism surjectivity property. Therefore, Theorem 11.6 implies that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$.
@@
Recall from the assumptions that $\mathsf{M}_{1}$ is a pp expansion of $\mathsf{K}$ and that $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$. Therefore, $\mathsf{M}_{2}$ is also a pp expansion of $\mathsf{K}$ by Theorem 10.6. This fact will be used repeatedly in the proof.
@@
We will prove that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\in\mathsf{M}_{1}$. Recall from the assumptions that $\mathsf{K}$ is a universal class. Therefore, so is its pp expansion $\mathsf{M}_{1}$ by Theorem 10.3(i). Together with the assumption that $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$, this allows us to apply Proposition 9.6, obtaining $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\in\mathsf{M}_{1}$, as desired. Together with the above display and the assumption that $\{t_{i}:i\in I\}$ interpolates $f$ in $\mathsf{M}_{1}$, this implies that there exists $i\in I$ such that $f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}(a_{1},\dots,a_{n})=t_{i}^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}}(a_{1},\dots,a_{n})=t_{i}^{\bm{A}}(a_{1},\dots,a_{n})$.
(ii): Assume that $\mathsf{M}_{2}$ is a Beth companion of $\mathsf{M}_{1}$. We will prove that $\mathsf{M}_{2}$ is a Beth companion of $\mathsf{K}$ as well. As in the previous case, it suffices to show that every implicit operation of $\mathsf{K}$ is interpolated in $\mathsf{M}_{2}$ by a family of terms of $\mathsf{M}_{2}$. Accordingly, consider an $n$-ary $f\in\mathsf{imp}(\mathsf{K})$. By Proposition 10.26 there exists $g\in\mathsf{imp}(\mathsf{M}_{1})$ such that $g^{\bm{A}}=f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}}$ for each $\bm{A}\in\mathsf{M}_{1}$. Observe that $g$ is interpolated in $\mathsf{M}_{2}$ by a family $\{t_{i}:i\in I\}$ of terms of $\mathsf{M}_{2}$ because $\mathsf{M}_{2}$ is a Beth companion of $\mathsf{M}_{1}$. We will prove that the same family interpolates $f$ in $\mathsf{M}_{2}$. To this end, consider $\bm{A}\in\mathsf{M}_{2}$ and $a_{1},\dots,a_{n}\in A$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}})$. Since $\mathsf{K}$ is a universal class, Theorem 10.3(i) implies that $\mathsf{M}_{1}$ is a universal class as well. As $\bm{A}\in\mathsf{M}_{2}$ and $\mathsf{M}_{2}$ is a pp expansion of $\mathsf{M}_{1}$ we have $\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}\in\mathsf{M}_{1}$ by Proposition 9.6. Furthermore, $(\bm{A}{\upharpoonright}_{{L}_{\mathsf{M}_{1}}}){\upharpoonright}_{{L}_{\mathsf{K}}}=\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$ because ${L}_{\mathsf{K}}\subseteq{L}_{\mathsf{M}_{1}}$. Therefore,
@@
Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, we may assume that it is induced by some $\mathcal{G}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and $\mathcal{G}$-expansion ${L}_{\mathcal{G}}$ of ${L}_{\mathsf{K}}$. As $\mathcal{F}=\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ by assumption, we have $\mathcal{G}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})=\mathcal{F}$. Then there exists an $\mathcal{F}$-expansion ${L}_{\mathcal{F}}^{\prime}$ of the form ${L}_{\mathcal{G}}\cup\{g_{f}:f\in\mathcal{F}-\mathcal{G}\}$. Since $\mathcal{G}\subseteq\mathcal{F}$ and ${L}_{\mathcal{G}}\subseteq{L}_{\mathcal{F}}^{\prime}$, Theorem 10.7 implies that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}^{\prime}])$ is a pp expansion of $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{G}}])$. Together with the assumption that $\mathsf{M}$ is a Beth companion of $\mathsf{K}$, this allows us to apply Theorem 11.4(i), obtaining that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}^{\prime}])$ is a Beth companion of $\mathsf{K}$ as well. Lastly, the definition of ${L}_{\mathcal{F}}^{\prime}$ guarantees that the classes $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ and $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}^{\prime}])$ are term equivalent. ∎
@@
Applying (iv) to $g\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ yields a set $\{t_{i}:i\in I\}$ of terms of $\mathsf{M}$ that interpolates $g$ in $\mathsf{M}$. We will show that $\{t_{i}:i\in I\}$ interpolates $f$ as well. To this end, consider $\bm{A}\in\mathsf{M}$ and $a_{1},\dots,a_{n}\in A$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$. Since $\bm{A}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, we have $\bm{A}\leqslant\bm{B}$ for some $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$. As $f\in\mathsf{imp}(\mathsf{M})$ and $\bm{A},\bm{B}\in\mathsf{M}$ are such that $\bm{A}\leqslant\bm{B}$, from Proposition 8.1 and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{A}})$ it follows that
@@
It only remains to prove Theorem 11.7. The proof hinges on the next observation.
@@
Let $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ be a pair of Beth companions of a quasivariety $\mathsf{K}$. We will prove that $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are faithfully term equivalent relative to $\mathsf{K}$. For $i=1,2$ let $\mathcal{F}_{i}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ be such that $\mathsf{M}_{i}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{i}}])$. We may assume that
@@
Then let $T_{i}$ be the set of terms of ${L}_{\mathcal{F}_{i}}$ with variables in $\{x_{n}:n\in\mathbb{N}\}$. Since $\mathsf{K}$ is a quasivariety, Theorem 11.6 implies that for every $n$-ary $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$ and $i=1,2$ there exists an $n$-ary term $F_{i}(f)\in T_{i}$ that interpolates $f$ in $\mathsf{M}_{i}$. Therefore, for every $n$-ary $f\in\mathsf{imp}_{\textsc{pp}}(\mathsf{K})$, $\bm{A}\in\mathsf{M}_{i}$ with $i=1,2$, and $a_{1},\dots,a_{n}\in A$,
@@
for each $f_{1}\in\mathcal{F}_{1}$ and $f_{2}\in\mathcal{F}_{2}$. We proceed to show that $\tau$ and $\rho$ witness that $\mathsf{M}_{1}$ and $\mathsf{M}_{2}$ are faithfully term equivalent relative to $\mathsf{K}$. Let $\mathcal{F}=\mathcal{F}_{1}\cup\mathcal{F}_{2}$ and let ${L}_{\mathcal{F}}$ be an $\mathcal{F}$-expansion of ${L}_{\mathsf{K}}$ such that ${L}_{\mathcal{F}_{1}},{L}_{\mathcal{F}_{2}}\subseteq{L}_{\mathcal{F}}$. Recall that $\mathcal{F}=\mathcal{F}_{2}\cup\mathcal{F}_{2}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ by assumption. Therefore, Theorem 10.7 implies that $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of both $\mathsf{M}_{1}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{1}}])$ and $\mathsf{M}_{2}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}_{2}}])$. To verify conditions (i) and (iii) in the definition of faithful term equivalence, consider $\bm{A}\in\mathsf{M}_{1}$. We need to prove that $\rho(\bm{A})\in\mathsf{M}_{2}$ and $\tau(\rho(\bm{A}))=\bm{A}$. Since $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{M}_{1}$, Proposition 10.2 implies that $\bm{A}$ is an ${L}_{\mathcal{F}_{1}}$-subreduct of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. As every member of $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a subalgebra of a member of $\mathsf{K}[{L}_{\mathcal{F}}]$, it follows that there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{1}}}$. Since $\mathsf{M}_{2}$ is a universal class by Theorem 10.3(i) and $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is also a pp expansion of $\mathsf{M}_{2}$, Proposition 10.2 implies that $\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{2}}}\in\mathsf{M}_{2}$.
We now show that $\rho(\bm{A})\in\mathsf{M}_{2}$. Observe that $\mathsf{M}_{2}$ is closed under subalgebras because it is a pp expansion of $\mathsf{K}$. Hence, it is sufficient to prove $\rho(\bm{A})\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{2}}}$ because $\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{2}}}\in\mathsf{M}_{2}$. Since $\bm{A}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathcal{F}_{1}}}$, from Remark 11.12(iii) it follows that
@@
Next we prove that $\tau(\rho(\bm{A}))=\bm{A}$. By Remark 11.12(iii) we obtain